1/3

Ирина Саверьевна Алексеева , Ирина Саверьевна Алексеева , Ирина Саверьевна Алексеева , Ирина Саверьевна Алексеева ,

Ode a Samarcanda Ode to Samarkand Ode für Samarkand

Oда Самарканду

tradukita de Carlo Minnaja tradukita de Malcolm Jones tradukita de Manfred Retzlaff

Argiope-informo: La aspekto de tiu ĉi

versio estas barita.

Supozeble la aŭtoro decidis tion.

I sing of you, oh, Samar-

kand’! Your wondrous features

win my praises. And your dear heart - the

Registan - seems young to me and

never ages.

Ich singe dir, o Samar-

kand, Dies Lied. Du bist so jung geblieben. Du bist in aller Welt

bekannt. Wer dich gesehn, der muss dich lieben.

Пою  тебя,  о,

Самарканд! Твой чудный облик восхваляю. И твоё сердце - Регистан - Я вечно юным называю.

Its steady beating Ulugh-

bek could sense beneath the

constellations. Oh, Samarkand! Immor-

tal pick of mighty Tamerlan’s

creations.

Ein Seher hat dich einst geschaut Im Traum vor langer, langer Zeit, Und Tamerlan hat dich gebaut In deiner ganzen Herrlichkeit.

Его биенье Улугбек Предчувствовал

под небом звёздным, О, Самарканд! Твой долгий век Величием Тимура создан!

Madrasses are your sa-

cred schools, which cultivate the arts

and learning. Their beauty overcomes

time’s rules by centuries into mo-

ments turning.

Und deine heil’gen Hallen, Hort Der Kunst, der Wissenschaft und Lehre,

Jahrhunderte schon stehn sie dort. Als ob kaum Zeit vergangen wäre.

Твои святые медресе - Приют искусств и просвещенья, Они стоят во всей

красе Столетья превратив в мгновенья.

Here the great Navoi’s

verse, so full of love and strong

emotion, is cherished by your po-

pulace, made happy by its inspi-

ration.

Hier hat der Dichter

Navoí Von Liebe und von Kraft gesungen, Und seine Lieder sind

noch nie In diesem Lande ganz verklungen.

Здесь Алишера

Навои Поэзию - любовь и

силу - Впитали жители

твои Она их счастьем вдохновила.

…  …  …  …

2/3

The domes of minarets

stand proud - like day-time stars in cle-

ar sky shining: so pure the blue, without

a cloud and fine the spires in art

combining.

Der Minarette Kup-

peln hier, Sie leuchten tags wie nachts die Sterne. Mit ihrer Ornamente

Zier Sie grüßen jeden schon von ferne.

Здесь минаретов купола - Дневные звёзды в небе ясном - Так чистая лазурь

светла И вязь орнаментов прекрасна.

Oh, Samarkand! You’re

wise through age, with mosques and chur-

ches closely neighboured.

For those who love you,

the world’s stage has no town better nor

more favoured.

O Samarkand, der Weisheit Stadt! Hier stehen Kirchen und Moscheen, Und wer dich liebt, der sagt, er hat Nichts Schön’res auf der Welt gesehen.

О, Самарканд! Ты - мудрость лет, Здесь рядом церкви и мечети… В тебя влюблённым лучше нет И краше города на

свете!

Oh, Samarkand! Live long

and thrive, be glorified by each good

action! And while eternity’s your

prize, be here our lasting bene-

faction.

O Samarkand, blüh’ und gedeih’, Dein Ruhm sich überall verbreite, Und ewig unvergänglich sei, Auch in der Zukunft uns begleite!

О, Самарканд! Живи, цвети И славься добрыми делами!

Навстречу вечности

лети И вечно оставайся с

нами.

…  …  …  …

3/3

Traduko de la Rusa poemo “Oда Самарканду” de Ирина Саверьевна Алексеева en itala de Carlo Minnaja (∗1940-03-19).

Arg-860-2206 (2014-11-13 12:31:20)

Tiu ĉi poem-italigo aperis en 2013 en la libro “Oда Самарканду”. Ода Самарканду: пер. на 62 языка / И. С. Алексеева ; ред.- сост.: А. Ионесов .

– Москва. : Изд-во журн. “Юность”, 2013 . – 91 с. : ил., 2 л. ил. Pri la tradukinto vidu la retejon http: // it. wikipedia. org/ wiki/ Carlo_ Minnaja .

Traduko de la Rusa poemo “Oда Самарканду” de Ирина Саверьевна Алексеева en la Anglan de Malcolm Jones.

Arg-860-2037 (2014-10-16 20:52:12)

Tiu ĉi poem-angligo aperis en 2013 en libro “Oда Самарканду” (= Odo por Samarkando). Sinjoro Malcolm Jones, kiu angligis la poemon, estas brita esperantisto. Lia adreso troviĝas en la UEA-jarlibro de la jaro 2014 sur paĝo 126.

Traduko de la Rusa poemo “Oда Самарканду” de Ирина Саверьевна Алексеева en la Germanan de Manfred Retzlaff (∗1938-11-04).

Arg-860-1735 (2013-04-03 23:31:01)

Verkinto de tiu ĉi Rusa poemo estas Ирина Саверьевна Алексеева .

Arg-860-1733 (2014-05-15 12:22:00)

Tiu ĉi poemo aperis en 2013 en la libro “Oда Самарканду”. Ода Самарканду: пер. на 62 языка / И. С. Алексеева ; ред.- сост.: А. Ионесов .

– Москва. : Изд-во журн. “Юность”, 2013 . – 91 с. : ил., 2 л. ил.